OpenStudy (anonymous):

Can anyone help me? Find the first 4 partial sums. The answers can be left in exact form or written as 4-place decimals. cos({n*pi}/4) for n=0 to infinity

OpenStudy (anonymous):

for \(n=0\) you get \(\cos(0)=1\)

OpenStudy (anonymous):

for \(n=1\) you get \(\cos(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\) so second partial sum is \[1+\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

for \(n=2\) you get \(\cos(\frac{\pi}{2})=0\) so third partial sum is the same as the second

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

Awesome Thank you!

OpenStudy (anonymous):

yw